# Scaling Up Active Testing to Large Language Models

### v1

@@ 4 Estimating Risk-Estimation Error
Aside from its scalability, active testing poses a practical challenge in assessing the quality of the risk estimates it produces. In particular, while the accuracy of simple Monte Carlo error estimates can easily be assessed through calculating the standard error across datapoints, previous work on active testing has not provided any comparable single run error estimator. Given practical deployment of active testing by definition entails only a single run, this leaves practitioners unaware of how well their implementation is working, making it harder to justify active testing’s use.
We address this by proposing a novel estimator of the mean squared error (over possible sequences of acquired indices, $i_{1:K}\sim q(i_{1:K})$, where $K\leq M$) of $\hat{R}_{\mathrm{LURE}}$. For a generic risk estimator, $\hat{R}(i_{1:K})$, we can decompose the mean squared error into bias and variance:
[…]
Letting $L=\{v_{m}\ell(f(x_{i_{m}}),y_{i_{m}})\}_{m=1}^{K}$ denote the sequence of precomputed reweighted losses (as used in Equation˜1), a bootstrap estimator [7] of the risk, $R$, is $\hat{R}_{\mathrm{boot}}=\frac{1}{K}\sum_{m=1}^{K}L_{j_{m}}$, where $j_{m}\sim\mathrm{Uniform}(1,2,\ldots,K)$. Now if we have $B$ bootstrap estimates of the risk, $(\hat{R}_{\mathrm{boot}}^{b})_{b=1}^{B}$, we can form a bootstrap estimator of $\textsc{Var}(\hat{R}_{\mathrm{LURE}})=\mathrm{MSE}(\hat{R}_{\mathrm{LURE}})$:
$\widehat{\textsc{Var}}_{\mathrm{boot}}(\hat{R}_{\mathrm{LURE}})=\frac{1}{B-1}\sum_{b=1}^{B}(\hat{R}_{\mathrm{boot}}^{b}-\bar{R}_{\mathrm{boot}})^{2},$ (6)
where $\bar{R}_{\mathrm{boot}}$ is the mean of the bootstrap estimates, namely $\bar{R}_{\mathrm{boot}}=\frac{1}{B}\sum_{b=1}^{B}\hat{R}_{\mathrm{boot}}^{b}$. We note that $\hat{R}_{\mathrm{boot}}$ is not a standard bootstrap estimator because the reweighted losses, $L_{j}$, are not independent and identically distributed (they are the result of active data acquisition), and the bootstrap resampling process does not capture the dependencies between the losses. As a result, $\hat{R}_{\mathrm{boot}}$ lacks theoretical convergence guarantees. Nevertheless, in Section˜5.9 we find it to be reliably accurate for our experiments, and thus that it provides useful estimates.
@@ 5 Experiments
We now seek to empirically assess our proposed active testing approach for LLMs, as well as our proposed risk-error-estimation method (Section˜4). We provide implementation details in Appendix˜A and code for running our experiments at anonymous.4open.science/r/scaling-up-active-testing-0.
@@ 5.1 Setup
We focus our evaluation on the following five well-established text classification tasks: Stanford Sentiment Treebank 2 (SST-2; 31), Subjectivity (Subj; 26), Financial Phrase Bank (FPB; 22), Hate Speech (HS; 6), and Massive Multitask Language Understanding (MMLU; 36).
The models we use are the 7B and 70B versions of Llama-2 [34] and Gemma-3 4B [33]. We use these models with either zero- or few-shot prompting. For Llama, we denote the four configurations as 7B-zero, 70B-zero, 7B-few and 70B-few; for Gemma, we use Gemma-3 4B-zero and Gemma-3 4B-few. We always include appropriate task prompts such as “Classify the sentiment of the following sentence as positive or negative”. When we use a few-shot surrogate with $n$ labelled in-context examples to evaluate a zero-shot target, we ensure a fair comparison with uniform sampling by comparing our results using $x$ acquired labels to random sampling results using $x+n$ acquired labels.
Figure 2: Cheap surrogate models support effective data acquisition for active testing. We compare uniform-random testing to active testing (LURE) across four datasets and four target-model configurations (7B-zero, 70B-zero, 7B-few and 70B-few). To guide active data acquisition we use a surrogate (Llama2-7B/-70B in (a), Gemma3-4B in (b)) that we train using a single step of in-context learning and then fix. This stripped-back surrogate-model training, along with the use of small surrogate models relative to the target models, allows us to drastically reduce the computational cost of active testing while achieving strong performance.
We use a logarithmic loss, as discussed in Sections˜2 and 3. Like in past work [16, 17], we measure the performance of a testing method using the median squared error of its risk estimate across random seeds, where the ground-truth risk is computed using a held-out test set.
@@ 5.2 Fixed surrogate models trained by in-context learning support effective data acquisition
To show the benefits of active testing for risk estimation more clearly, we plot the active risk-estimation error as a ratio of the uniform-random risk-estimation error, which we refer to as the relative error. We find that active testing performs very well relative to the standard uniform-random testing in the vast majority of settings (Figure˜2(a)), with risk-estimation error lowered by 32% on average (median over all the experimental variables listed above, as well as over possible label budgets between 1 and 400). This is an exciting result: a practically straightforward technique allows us to improve on standard practice for LLM evaluation with minimal computational overhead. Contrasting with the overall positive result, LURE-70B performs poorly in evaluating the 70B-few model on the SST-2 dataset. We explore this failure case in Section˜5.8 and show it is due to a large number of incorrect labels in the dataset.
Figure 3: Models can both be targets and surrogates in active testing. Although Gemma-3 is smaller than the Llama family, its capabilities are comparable. Llama models—including the 7B variant—can effectively evaluate it, demonstrating that older models can serve as reliable surrogates for newer ones.
5.3 Small surrogate models can be used to evaluate much larger target models
A key additional result to highlight in Figure˜2(a) is that LURE-7B works well not just for evaluating a target model of the same size, but also for evaluating a target model ten times larger. In fact, we show that we can reduce surrogate costs even further by using Gemma-3 4B (Figure˜2(b), bottom row) and Phi-2 (Figure˜C.1 in the Appendix) to evaluate Llama models, reducing computational costs by using even smaller surrogates. Conversely, Llama models effectively evaluate Gemma-3, whose capabilities rival those of larger Llama models on these datasets (Tables C.7 and C.8). As discussed in Section˜3.2, this ability to use a small surrogate model—combined with the fact that fixing the surrogate model allows us to only compute one forward pass on the pool set—is a useful way of reducing costs.
@@ 5.4 Using stripped-back surrogate models favours sampling-based active testing
Next we evaluate our claim in Section˜3.1 that sampling-based active testing is preferable over interpolation-based active testing given that the surrogate model is fixed. We do this by comparing our approach with ASE in evaluating the 7B-few and 70B-few models. Our results show ASE produces strong results on the HS dataset with the 70B-few model, but otherwise significantly underperforms (Figure˜4). This is consistent with our explanation that ASE is more sensitive to the quality of the surrogate model, and that computational constraints on the surrogate model make it less practically useful. Note that the ASE risk estimator is constant if the surrogate model is fixed, as it is here.
Figure 4: Our sampling-based active testing (LURE) approach works more reliably than interpolation-based active testing (ASE) when using cheaply constructed surrogate models. In ASE, the surrogate model not only guides data acquisition but also much more directly affects the risk estimate: the labels used to compute the expected loss of the target model are simulated from the surrogate model.
@@
5.5 Good data acquisition is possible without computing predictions with the target model
@@
Now we explore our suggestion in Section˜3.3 that the surrogate model can be used to approximate the target model to completely avoid the need for any target model evaluations during data acquisition. Since the loss is logarithmic, this equates to using the predictive entropy of the surrogate model as the acquisition function. We assess this approach as applied to evaluating the 70B-few model with LURE-7B. We find using the surrogate model to approximate the target model during data acquisition is surprisingly effective (Figure˜5(a)). This finding is very promising: it suggests all of the cost-reduction measures we proposed in Section˜3 are compatible with strong active-testing performance.
Figure 5: (5(a)) Approximating the target model with the surrogate model during data acquisition (here this means acquiring based on the predictive entropy of the surrogate model) reduces the need to compute predictions with the target model while still providing good data acquisition. (5(b)) Active testing with a label-aware acquisition function (here the negative log likelihood of the surrogate model) enables dataset curation tailored to a target model of interest. In the curation problem setting we have access to labels for all inputs in the pool set and the goal is to select a subset of them to reduce the number of target-model predictions required to estimate its risk.
Figure 6: Active testing on the challenging MMLU task remains efficient in most cases, with graceful failure in hardest settings. Notably, Gemma-3 4B is an efficient surrogate, despite belonging to a different model family.
5.6 Active testing for dataset curation is viable
@@
Here we consider the alternative problem setting discussed in Section˜3.4, where the inputs in the pool set are already labelled but it is too expensive to compute target-model predictions on all of them, with a possible solution being to curate a subset using a cheap surrogate model. More concretely we run active testing using the negative log likelihood as our label-aware acquisition function (Equation˜4). We find that our approach improves over uniform-random subsampling on three datasets out of four, and roughly matches on FPB (Figure˜5(b)). These results—along with those for the predictive-entropy function we considered in Section˜5.5—suggest that active testing could be a useful tool for dataset curation, reducing computational costs when using existing evaluation datasets. The practical benefits of this could be significant: Liang et al, [21] required almost 20,000 hours of (Nvidia A100) GPU hours to evaluate 30 models on their HELM benchmark.
@@
Figure 7: The failure of LURE-70B in efficiently evaluating the 70B-few model on the SST-2 dataset (left) is resolved by filtering incorrectly labelled datapoints from the pool set (middle, right). We filter points with a high loss (NLL) and manually confirm that most labels for these points are incorrect (see main text).
5.7 Active evaluation on harder tasks: the MMLU dataset
We next assess the robustness of active testing on the more challenging MMLU dataset [36], where model—and thus surrogate—performance degrades substantially (Tables C.7 and C.8), making data acquisition harder. Despite this, active testing continues to outperform uniform sampling, with graceful failures in the hardest evaluation setting, 70B-few in (Figure˜6). As task difficulty increases, weaker surrogates become insufficient but fail gracefully, as discussed in Section˜C.2. Interestingly, Gemma-3 4B is particularly efficient to evaluate Llama models, highlighting the benefits of diversity when choosing a surrogate from a different model family than the target for cross-entropy sampling. These results show that even small, sufficiently capable models can enable effective active testing.
5.8 Incorrect labels can cause problems
@@
Now we investigate the failure of LURE-70B to effectively evaluate the 70B-few model on the SST-2 dataset (Section˜5.2). We construct two modified versions of the SST-2 pool set: one filtered to remove inputs for which 70B-few’s negative log likelihood (NLL) is greater than 5, and one filtered with a NLL threshold of 3. Then we re-run active testing with these modified pool sets.
@@
We show that the failure case is reduced by the weaker filtering and completely resolved by the stronger filtering (Figure˜7). To understand why, we inspect the filtered-out data and find it is often mislabelled. We find 75% of the 45 examples with NLL greater than 5 are mislabelled, and 67% of the 136 examples with NLL greater than 3. This is a useful demonstration of how active testing can fail when the source of labels is unreliable. Interestingly it can also be seen as a case for using less powerful surrogate models to guide data acquisition: LURE-7B worked well in the exact setting where LURE-70B failed. It is also worth noting that in a dataset-curation setting (Section˜5.6) we have access to all the test labels and therefore have the option to directly filter out high-NLL examples.
@@
Figure 8: Our bootstrap estimator of the risk-estimation error in active testing provides a useful indicator of performance within a given data-acquisition rollout. High coverage probabilities mean the confidence interval is very likely to contain the true risk-estimation error.
@@ 5.9 Bootstrap estimation of risk-estimation error provides an accurate performance indicator
Having assessed our proposed changes to active testing, we turn to our proposed method for estimating the mean squared error of the LURE risk estimator (Section˜4). We run active testing $T=100$ times, generating 100 sequences of acquired indices, $(i_{1:M})_{j=1}^{100}$. Then for $K\in\{1,2,\ldots,M\}$ we perform four steps. First, we compute a ground-truth mean-squared error, $\mathrm{MSE}_{K}$, using Equation˜5 with $q(i_{1:K})$ defined as a uniform distribution over the $K$ index sequences acquired so far. Second, we compute a boostrap estimate of the MSE, $\widehat{\mathrm{MSE}}_{K}$, using Equation˜6 with $B=1000$ bootstrap risk estimates. Third, we compute the MSE-estimation error, $(\mathrm{MSE}_{K}-\widehat{\mathrm{MSE}}_{K})^{2}$. Fourth, we construct an approximate 95% confidence interval, $\widehat{\mathrm{MSE}}_{K}\pm 2\hat{\sigma}$, where $\hat{\sigma}$ is the estimated standard deviation.
Our results show MSE-estimation error converging to low values for most runs, after an initial period ($K<100$) of relatively high error (Figure˜8). In addition to this, our confidence intervals provide high levels of coverage: for a strong majority of runs we see coverage probabilities of around 94% for $K\geq 100$, suggesting a practitioner can expect their computed confidence interval to contain the true mean squared error approximately 94% of the time. While there are cases where MSE estimation fails, the overall performance is promising, suggesting our proposed method could be a useful practical performance indicator, helping justify real-world deployment of active testing.

### v2

@@ 4 Estimating risk-estimation error
Aside from its scalability, active testing poses a practical challenge in assessing the quality of the risk estimates it produces. In research we can construct repeatable simulations of active testing and record ground-truth risk-estimation errors across repeat runs. But practical deployment of active testing involves only a single run, without knowledge of the true risk. This leaves practitioners unaware of how well their implementation is working, making it harder to justify using active testing.
We propose addressing this using a novel estimator of the mean squared error (over possible sequences of acquired indices, $i_{1:K}\sim q(i_{1:K})$, where $K\leq M$) of $\hat{R}_{\mathrm{LURE}}$. For a generic risk estimator, $\hat{R}(i_{1:K})$, we can decompose the mean squared error into squared bias and variance:
$\mathrm{MSE}(\hat{R})=\mathbb{E}_{q(i_{1:K})}[(\hat{R}(i_{1:K})-R)^{2}]=\underbrace{(\mathbb{E}_{q(i_{1:K})}[\hat{R}(i_{1:K})]-R)^{2}}_{\mathrm{Bias}(\hat{R})^{2}}+\underbrace{\mathbb{V}_{q(i_{1:K})}[\hat{R}(i_{1:K})]}_{\mathrm{Var}(\hat{R})}$. (5)
Since $\mathrm{Bias}(\hat{R}_{\mathrm{LURE}})=0$, our task is to estimate $\mathrm{Var}(\hat{R}_{\mathrm{LURE}})=\mathrm{MSE}(\hat{R}_{\mathrm{LURE}})$ from a single sequence of acquired indices, $i_{1:K}$. Letting $L=(v_{m}\ell(f(x_{i_{m}}),y_{i_{m}}))_{m=1}^{K}$ denote the first $K$ reweighted losses used in Equation 1, a bootstrap estimator (efron1979bootstrap) of the risk, $R$, is
$\hat{R}_{\mathrm{boot}}=\frac{1}{K}\sum_{m=1}^{K}L_{j_{m}},\quad j_{m}\sim\mathrm{Uniform}(1,2,\ldots,K).$
If we have $B$ bootstrap estimates, $(\hat{R}_{\mathrm{boot}}^{b})_{b=1}^{B}$, we can form a bootstrap estimator of the variance:
$\widehat{\mathrm{Var}}_{\mathrm{boot}}(\hat{R}_{\mathrm{LURE}})=\frac{1}{B-1}\sum_{b=1}^{B}(\hat{R}_{\mathrm{boot}}^{b}-\bar{R}_{\mathrm{boot}})^{2},\quad\bar{R}_{\mathrm{boot}}=\frac{1}{B}\sum_{b=1}^{B}\hat{R}_{\mathrm{boot}}^{b}.$ (6)
It is important to note that $\hat{R}_{\mathrm{boot}}$ is not a standard bootstrap estimator because the reweighted losses, $L_{j}$, are not independent and identically distributed (they are the result of active data acquisition), and the bootstrap resampling process does not capture the dependencies between the losses. As a result, the estimator lacks theoretical convergence guarantees. Nevertheless we find it to be reliably accurate in our experiments (Section 5.9), suggesting it could be useful in practice.
@@ 5 Experiments
We now seek to empirically assess our proposed approach for active testing of LLMs (Section 3), as well as our proposed risk-error-estimation method (Section 4). We provide implementation details in Appendix B and code at github.com/gabrielleberrada/scaling-up-active-testing.
@@ 5.1 Setup
We use five text-classification datasets: Stanford Sentiment Treebank 2 (SST-2; socher2013recursive), Subjectivity (Subj; pang2004sentimental), Financial Phrase Bank (FPB; malo2013good), Hate Speech (HS; degibert2018hate) and Massive Multitask Language Understanding (MMLU; hendrycks2021measuring). We mainly focus on the first four, which were previously used by kossen2024context, then use MMLU to explore a problem that is more challenging for the models used in our experiments.
The models we use span three model families and have parameter counts ranging from 1 to 70 billion. In our main experiments we use the 7B and 70B versions of Llama 2 (touvron2023llama) along with Gemma3-4B (kamath2025gemma); in additional experiments we use Phi-2 (javaheripi2023phi) and Gemma3-1B. We use these models with either zero- or few-shot prompting, with both including appropriate task prompting such as “Classify the sentiment of the following sentence as positive or negative”. We denote our four Llama 2 model configurations as 7B-zero, 70B-zero, 7B-few and 70B-few, and our two Gemma 3 model configurations as Gemma3-4B-zero and Gemma3-4B-few.
(a) Llama 2 target models and Llama 2 surrogate models
(b) Llama 2 target models and Gemma3-4B surrogate model
Figure 2: Cheap surrogate models support effective data acquisition for active testing. We compare uniform-random testing to active testing (LURE) across four datasets. To guide active data acquisition we use a surrogate model that we train using a single step of in-context learning and then keep fixed. This stripped-back surrogate-model training, along with the use of small surrogate models relative to the target models, allows us to drastically reduce the computational cost of active testing while achieving strong performance.
We compare active testing to uniform-random testing, a standard approach to model evaluation. Acquisition methods from the active-learning literature (settles2012active) are not tailored for risk estimation (kossen2021active) and so would have limited utility as baseline methods here.
We use a logarithmic loss, as discussed in Sections 2 and 3. Like in past work (kossen2021active; kossen2022active), we measure the performance of a testing method using the median squared error of its risk estimate across random-number seeds, where the ground-truth risk is computed using a held-out test set. Alongside plots of the absolute value of this risk-estimation error, we present plots of relative error, namely the active risk-estimation error divided by the uniform-random risk-estimation error. Visualising performance on this relative scale lets us see the benefits of active testing more clearly. When we use a few-shot surrogate model with $n$ labelled in-context examples to evaluate a zero-shot target model, we ensure a fair comparison to uniform-random testing by comparing our results for $k$ acquired labels to uniform-random results for $n+k$ acquired labels.
@@ 5.2 Fixed surrogate models trained by in-context learning support effective data acquisition
We find that active testing performs very well relative to the standard uniform-random testing in the vast majority of cases (Figure 2(a)), with risk-estimation error lowered by 32% on average (median over all the experimental variables listed above, as well as over possible label budgets between 1 and 400). This is an exciting result: a practically straightforward technique allows us to improve on standard practice for LLM evaluation with minimal computational overhead.
Contrasting with the overall positive result in Figure 2(a), LURE-70B performs poorly in evaluating the 70B-few model on the SST-2 dataset. We explore this failure case in Section 5.8 and show it is due to a large number of incorrect labels in the dataset. In Section C.5 we present an additional experiment that compares dynamic surrogate models to fixed ones: for the scenarios we study, dynamic surrogate models provide no accuracy benefits but incur prohibitive computational costs.
Figure 3: Older Llama 2 models are useful surrogates for active testing of newer Gemma 3 target models.
Figure 4: Our sampling-based active testing (LURE) approach works more reliably than interpolation-based active testing (ASE) when using cheaply constructed surrogate models. In ASE the surrogate model not only guides data acquisition but also much more directly affects the risk estimate: the labels used to compute the expected loss of the target model are simulated from the surrogate model.
5.3 Small surrogate models can be used to evaluate larger target models
A key additional result to highlight in Figure 2(a) is that LURE-7B works well not just for evaluating a target model of the same size, but also for evaluating a target model ten times larger. In fact, we show that we can reduce surrogate-model costs even further by using Gemma3-4B (Figure 2(b)) and Phi-2 (Figure 10(a)) to evaluate Llama 2 models, reducing computational costs by using even smaller surrogates. Meanwhile, older Llama 2 models effectively evaluate newer Gemma 3 models (Figure 3), which perform strongly on these datasets (Tables 3 and 2). As discussed in Section 3.2, this ability to use a small surrogate model—combined with the fact that fixing the surrogate model allows us to only compute one forward pass on the pool set—is a useful way of reducing costs.
@@ 5.4 Using stripped-back surrogate models favours sampling-based active testing
Next we evaluate our claim in Section 3.1 that sampling-based active testing is preferable over interpolation-based active testing given that the surrogate model is fixed. We do this by comparing our approach with ASE in evaluating the 7B-few and 70B-few models. Our results show ASE produces strong results on the HS dataset with the 70B-few model, but otherwise significantly underperforms (Figure 4). This is consistent with our explanation that ASE is more sensitive to the quality of the surrogate model, and that computational constraints on the surrogate model make it less practically useful. Note that the ASE risk estimator is constant if the surrogate model is fixed, as it is here.
@@
5.5 Good data acquisition is possible without making predictions with the target model
@@
Now we explore our suggestion in Section 3.3 that the surrogate model can be used to approximate the target model, such that we do not require target-model predictions during data acquisition. For the logarithmic loss we are using, this equates to using the predictive entropy of the surrogate model as the acquisition function. We assess this approach as applied to evaluating the 70B-few model with LURE-7B, and find it is surprisingly effective (Figure 6). This is promising: it suggests all the cost-reduction measures proposed in Section 3 are compatible with strong active-testing performance.
Figure 5: Using the surrogate model to approximate the target model during data acquisition (here this means acquiring data based on the predictive entropy of the surrogate model) reduces the need for target-model predictions while still performing well.
Figure 6: Active testing with a label-aware acquisition function (here the negative log likelihood of the surrogate model) enables dataset curation (selecting a subset of already-labelled test data) tailored to a target model of interest, helping reduce computational costs.
5.6 Active testing for dataset curation works, suggesting scope to reduce computational costs
@@
Here we consider the alternative problem setting discussed in Section 3.4, where the inputs in the pool set are already labelled but it is too expensive to compute target-model predictions on all of them, with a possible solution being to curate a subset using a cheap surrogate model. More concretely we run active testing using the negative log likelihood as our label-aware acquisition function (Equation 4).
We find our approach improves over uniform-random subsampling on three datasets out of four (SST-2, Subj and HS) and performs slightly worse on FPB (Figure 6). These overall-positive results—along with those for the predictive-entropy function we considered in Section 5.5—suggest that active testing could be a useful tool for dataset curation, reducing computational costs when using existing evaluation datasets. The practical benefits of this could be significant: liang2023holistic required almost 20,000 hours of (Nvidia A100) GPU hours to evaluate 30 models on their HELM benchmark.
5.7 The benefit of active testing tends to hold in more challenging scenarios
Next we use the MMLU dataset to assess the robustness of active testing to an increase in task difficulty, which corresponds to worse surrogate-model performance (Tables 3 and 2) and thus poses a challenge for effective data acquisition. Our results show active testing continuing to outperform uniform-random testing for most combinations of surrogate and target models, with the only failure occurring when using 70B-few for both the surrogate and target models (Figure 7).
Figure 7: Active testing outperforms uniform-random testing in most cases on the harder MMLU dataset.
@@ 5.7 The benefit of active testing tends to hold in more challenging scenarios
Figure 8: The failure of LURE-70B in efficiently evaluating the 70B-few model on the SST-2 dataset (left) is resolved by filtering incorrectly labelled datapoints from the pool set (middle, right). Underlying this effect is the presence of incorrect labels in the original dataset: we find that inputs with high loss are often mislabelled.
5.8 Incorrect labels can cause problems for active testing
@@
Now we investigate the failure of LURE-70B to effectively evaluate the 70B-few model on the SST-2 dataset (Section 5.2). We construct two modified versions of the SST-2 pool set: one filtered to remove inputs for which 70B-few’s negative log likelihood (NLL) is greater than 5, and one filtered with a NLL threshold of 3. Then we re-run active testing with these modified pool sets.
@@
We show that the failure case is reduced by the weaker filtering and completely resolved by the stronger filtering (Figure 8). To understand why, we inspect the filtered-out data and find it is often mislabelled. We find 75% of the 45 examples with NLL greater than 5 are mislabelled, and 67% of the 136 examples with NLL greater than 3. This is a useful demonstration of how active testing can fail when the source of labels is unreliable. Interestingly it can also be seen as a case for using less powerful surrogate models to guide data acquisition: LURE-7B worked well in the exact case where LURE-70B failed. It is also worth noting that in a dataset-curation setting (Section 5.6) we have access to all the test labels and therefore have the option to directly filter out high-NLL examples.
@@
Figure 9: Our bootstrap estimator of the active-testing risk-estimation error provides a useful performance indicator. Coverage probability measures how often a confidence interval contains the true risk-estimation error.
@@ 5.9 Bootstrap estimation of risk-estimation error provides an accurate performance indicator
Having assessed our proposed active-testing method, we turn to our method for estimating the mean squared error of the LURE risk estimator (Section 4). We run active testing $T=100$ times, generating 100 sequences of acquired indices, $(i_{1:M})_{j=1}^{100}$. Then for $K\in\{1,2,\ldots,M\}$ we perform four steps:
  1. 1.
    Compute a ground-truth mean-squared error, $\mathrm{MSE}_{K}$, using Equation 5 with $q(i_{1:K})$ defined as a uniform distribution over the $K$ index sequences acquired so far.
  2. 2.
    Compute a bootstrap estimate of the MSE, $\widehat{\mathrm{MSE}}_{K}$, using Equation 6 with $B=1000$ estimates.
  3. 3.
    Compute the MSE-estimation error, $(\mathrm{MSE}_{K}-\widehat{\mathrm{MSE}}_{K})^{2}$.
  4. 4.
    Compute an approximate confidence interval, $\widehat{\mathrm{MSE}}_{K}\pm 2\hat{\sigma}$, using estimated standard deviation $\hat{\sigma}$.
Our results show MSE-estimation error converging to low values for most runs, after an initial period ($K<100$) of relatively high error (Figure 9). In addition to this, our confidence intervals provide high levels of coverage: for a strong majority of runs we see coverage probabilities of around 94% for $K\geq 100$, suggesting a practitioner can expect their computed confidence interval to contain the true mean squared error approximately 94% of the time. While there are cases where MSE estimation fails, the overall performance is promising, suggesting our proposed method could be a useful practical performance indicator, helping justify real-world deployment of active testing.